Solve the equation: 42 + 13 -2 = 1 Answer: =
Answer #1

Ans given eq. 4+3 -2 - 4 = थ+2 J4+13 99uase on both both Side 2 (1१५ 413) - (५+2 4 +13 21244.4+4 0-21-421 +4 -44-13 = --9 १२hence Once we Get 4=3